• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name W. Korsch

    Log entry time 00:00:05 on July20,2001

    Entry number 66796

    keyword=shift summary

    Comments:
    17:00 back from Hall. Tested effect of 'helicity glitch' on Phillips G&D generator.
    Conclusion: Phillips seems not be sensitive to glitches, if we run in toggle mode.
    17:04 Sweep Hall.
    17:30 ER2 not responding, Bodo killed Coda and rebooted everything.
    17:40 MCC called, beam back in 1 minute.
    Request: 5 muA, Raster off, BeO check
    17:42 Bodo killed Coda on Parity machine again and restarted it.
    17:45 Checked beam position on BeO and raster spot. Looks ok.
    17:55 Asked MCC to deliver 12 muA with raster on.
    17:58 MCC called problems with the raster.
    18:13 MCC called: have to reboot ROC for raster, should get beam soon.
    18:17 Still about 10-15 min before beam -> decided to do polarisation
    measurment.
    18:30 Raster problem fixed, polarisation measurements finished
    18:30 NMR: 48.4%
    18:35 EPR: 41.7%
    18:45 charge asymmetry large (~ -900 ppm), Piotr fixed it manually
    23:30 NMR measurement: 45.5%
    23:40 EPR measurement: 43.9%
    23:47 request beam l/2 plate in

    Run Summary:
    l/2 (beam) out, l/2(target) out
    18:40 2125/21125 large charge asymmetry
    18:50 12084/2126/21126
    20:21 12085/2127/21127
    21:50 12086/2128/21128
    ------------------------------
    l/2 (beam) in, l/2 (target) out
    23:50 12087/2129/21129